# Learning Shared Representations from Unpaired Data

### v1

@@ Abstract
Learning shared representations is a primary area of multimodal representation learning. The current approaches to achieve a shared embedding space rely heavily on paired samples from each modality, which are significantly harder to obtain than unpaired ones. In this work, we demonstrate that shared representations can be learned almost exclusively from unpaired data. Our arguments are grounded in the spectral embeddings of the random walk matrices constructed independently from each unimodal representation. Empirical results in computer vision and natural language processing domains support its potential, revealing the effectiveness of unpaired data in capturing meaningful cross-modal relations, demonstrating high capabilities in retrieval tasks, generation, arithmetics, zero-shot, and cross-domain classification. This work, to the best of our knowledge, is the first to demonstrate these capabilities almost exclusively from unpaired samples, giving rise to a cross-modal embedding that could be viewed as universal, i.e., independent of the specific modalities of the data. Our code IS publicly available at https://github.com/shaham-lab/SUE.
@@ 1 Introduction
The great success of unimodal models Devlin (2018); Liu (2019); Reimers (2019); Chen et al. (2020); Caron et al. (2021); Bardes et al. (2021) in the last decade, and the increasing demand for multimodal applications, have shifted large research attention into the cross-modal domain. Multimodal models present impressive performance on various cross-modal tasks, including image-text Radford et al. (2021); Li et al. (2022, 2023), speech-text Elizalde et al. (2023), video-text Xu et al. (2021) and medical-image-text Zhang et al. (2022a), to name a few.
A primary task in multimodal representation learning is the learning of shared representations, i.e., representation spaces to which instances from different modalities can be mapped, and in which such multimodal instances can be compared. Currently, models learning shared representations typically rely on vast amounts of paired data for training. For example, CLIP Radford et al. (2021) was trained on 400 million pairs of images and their captions (i.e., corresponding texts). This kind of supervision (i.e., pairing) is often costly, difficult, and might be even impossible to obtain in many domains and applications. In medical domains, for instance, obtaining a large number of samples is often challenging, as it typically depends on expert annotations. This supervision becomes both costly and time-intensive, or requires rare and specialized tests Hervella et al. (2019). Unlike paired data, unpaired data is significantly more accessible and available.
Figure 1: Empirical demonstration of universality. (a) Distances between corresponding random walks on image and text graphs from MSCOCO, compared to distances to randomly shuffled (non-matching) walks. Although constructed independently from unimodal features, corresponding walks exhibit significantly greater similarity. (b) Distances between paired and unpaired points in the shared space of aligned 2D spectral embeddings (SEs). Paired points are consistently closer, indicating that the independently learned SEs capture analogous structure across modalities (see App. A.5).
@@
Figure 2: Almost exclusively unpaired image retrieval. Retrieved images for custom captions on the MSCOCO dataset, trained with 100 pairs and 10k non-pairs. The retrieved images are highly similar semantically to the text queries, even though almost no pairs were available during training.
Specifically, modern pre-trained unimodal foundation models have a proven ability to represent semantics. For example, two given images have close embeddings if their semantic meaning is similar, and far otherwise. These similarities can be captured by a random walk process on the samples’ representations, suggesting that a random walk process defined on such unimodal representations should largely correspond to semantic similarity. Therefore, we can expect random walks defined on different unimodal representations that capture semantics well to be highly similar (see Fig. 1). This assumption is supported by substantial empirical evidence from recent works Huh et al. (2024); Fan et al. (2025).
Random walk processes are finite analogs of diffusion operators. Thereby, the similarity of random walks that are constructed from different, modality-dependent representations implies that the eigenfunctions of the corresponding diffusion operators will have universality properties (i.e., modality-invariance) Coifman (2020). Therefore, constructing a spectral embedding (SE) based on the leading eigenvectors of random walks, which are viewed as discrete approximations of the leading eigenfunctions of diffusion operators Belkin and Niyogi (2006); Shi (2015), enables us to take advantage of this concept even in the absence of paired samples. That is, these eigenfunctions can be learned independently from each modality using only unpaired data.
We demonstrate our arguments via a pipeline we call SUE: Spectral Universal Embedding, designed to uncover a universal embedding in a weakly-paired scenario, i.e., when large amounts of unpaired data are available alongside only a very small number of paired samples. To the best of our knowledge, this work is the first to learn shared representations from almost exclusively unpaired data. SUE consists of three steps (see Fig. 3): (i) SE, to extract universal features, (ii) CCA on very few samples, to align those features, and (iii) MMD-net, to provide additional non-linear alignment. We demonstrate SUE on vision and language domains and show that with only a small number of paired examples, unpaired data can be leveraged to learn rich multimodal representations. For instance, on the MSCOCO dataset, SUE improves retrieval performance by over 250% on average compared to contrastive learning trained with the same number of pairs. Fig. 2 demonstrates SUE’s ability to learn cross-modal properties with almost no paired data.
@@ 2 Related Work
Multimodal models have recently seen great improvements in rich-data modalities, such as text with image Radford et al. (2021); Li et al. (2022, 2023), speech Elizalde et al. (2023) and video Xu et al. (2021). Current state-of-the-art multimodal models rely on contrastive learning between corresponding pairs of unimodal representations. While these methods yield high-quality results, they require extremely large amounts of paired data, limiting their applicability to arbitrary domains. In contrast, the universal embedding offers an integration of any two domains, uncovering a shared space with only a few paired samples. Moreover, following recent research on the modality gap in multimodal learning Liang et al. (2022); Eslami and de Melo (2025); Schrodi et al. (2025), contrastive-based methods fail to be universal due to modalities residing on different manifolds. In contrast, a universal embedding should not exhibit this gap (see App. B for details).
@@
Unpaired samples have been used to replace certain applications of multimodal models without constructing shared representations, including medical image segmentation Yang et al. (2022); Valindria et al. (2018), image-to-speech translation Ma et al. (2019), image-to-image translation Zhu et al. (2017a, b); Choi et al. (2018), bilingual word embeddings Artetxe et al. (2017), and domain confusion Ganin and Lempitsky (2015); Ganin et al. (2016); Stark et al. (2020); Zhang et al. (2022b). Other methods incorporate unpaired data to complement additional signals, such as large paired datasets Nakada et al. (2023); Singh et al. (2022); Devillers et al. (2024); Girdhar et al. (2023) or classification labels Timilsina et al. (2024); Xi et al. (2024). In contrast, SUE uncovers a latent universal space between two unimodal domains that share semantic content, using unpaired data and only a small number of pairs.
Several works explore shared representation learning from limited to no supervision. Hoshen and Wolf (2018) proposed an unpaired variant of CCA, but as noted in their work, the method is unstable and requires many runs to yield satisfactory results, limiting its practicality. MACK Huang et al. (2022), while targeting image-text alignment, depends on a segmentation-to-text model trained with paired data, limiting its scope. Closer to our direction are Chen et al. (2023); Mo et al. (2023), which align modalities by referencing a set of paired examples. However, they overlook local structure in the unimodal embedding space - an important cue, since non-immediate neighbors are often uninformative in high-dimensional settings In contrast, SUE leverages local unimodal neighborhoods and a few weak pairs to construct a unified cross-modal representation.
@@
SE is known for its global structure preservation properties of unimodal manifolds Coifman et al. (2005); Nadler et al. (2006); Talmon et al. (2015). One perspective, for instance, is that the SE is the space in which the Euclidean distances represent the diffusion distances on the original manifold Coifman and Lafon (2006). The diffusion distance is a manifold-dependent metric, offering advantages over traditional metrics (e.g., Euclidean, cosine), especially in high-dimensional spaces Coifman et al. (2005). Recent works investigated the global structure preservation abilities of SE in the cross-modal scenario Eynard et al. (2015); Lederman and Talmon (2018); Yacobi et al. (2025). Specifically, they showed that joint SE extracts meaningful representations from each modality. The universal embedding is inspired by these findings, and the analysis we provide extends this idea. Particularly, we show that independent SEs capture similar semantic properties across different modalities, such as images and texts.
@@ 3 Preliminaries
SE Belkin and Niyogi (2003); Coifman and Lafon (2006) is a popular method, used across various domains McInnes et al. (2018); Defferrard et al. (2016); Zhang et al. (2021); Beaini et al. (2021); Dwivedi and Bresson (2020); Kreuzer et al. (2021); Kundu et al. (2004); Shepherd et al. (2007); Campbell et al. (2015); Zhu and Schlick (2021). In particular, let $\mathcal{X}\subseteq\mathbb{R}^{d}$ denote a collection of data points sampled from a manifold $\mathcal{M}$. Let $W\in\mathbb{R}^{n\times n}$ be a positive symmetric graph affinity matrix, with nodes corresponding to $\mathcal{X}$, and let $D$ be the corresponding diagonal degree matrix (i.e., $D_{ii}=\sum_{j=1}^{n}W_{ij}$). The random walk matrix is defined as $P=D^{-1}W$, and the random walk graph Laplacian as $L=I-P$, where $I$ is the identity matrix. The eigenvalues of $P$ can be sorted to satisfy $1=\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$ with corresponding eigenvectors $v_{1},v_{2},\dots,v_{n}$ Von Luxburg (2007).
@@
Given a Riemannian manifold $\mathcal{M}$ the Laplace-Beltrami operator $\Delta$ acts on functions $f:\mathcal{M}\to\mathbb{R}^{k}$ by $\Delta f=\text{div}(\nabla f)$, where div is the divergence and $\nabla$ the gradient. The diffusion operator (heat kenrnel) at time $t$ is then defined by $P_{t}=e^{-t\Delta}$. For more details please see Coifman and Lafon (2006). Importantly to this work, the Laplacian and random walk matrices converge to the Laplace-Beltrami and diffusion operators, respectively Coifman and Lafon (2006); Belkin and Niyogi (2006); Shi (2015).
@@
Recent works proposed parametric SE implementations, such as SpectralNet Shaham et al. (2018); Ben-Ari et al. (2025). These are deep-learning methods, designed to tackle the scalability and generalizability drawbacks of SE. Specifically, they learn a parametric map $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$, which minimizes the Rayleigh quotient
[…]
@@
MMD Gretton et al. (2006, 2012) is a distance measure between two probability distributions $p,q$. It is defined with respect to a function class $\mathcal{F}$ by
[…]
@@
[…]
@@ 4.1 Rationale
We formalize our assumption as follows. Let $\mathcal{M}$ be a latent, underlying semantic manifold, and let $f,g$ be two transformations, such that $f(\mathcal{M})$ and $g(\mathcal{M})$ represent the two modalities from which we observe samples. There is a body of work specifying conditions under which the spectral properties of $\mathcal{M}$ are preserved under $f,g$. For example, if $f,g$ have bounded distortion and bounded Ricci curvature, the corresponding eigenfunctions of the Laplace-Beltrami operator, as well as of the diffusion operator, on $f(\mathcal{M})$ and $g(\mathcal{M})$ are similar in the $L_{\infty}$ sense Bérard et al. (1994).
Intuitively, our assumption states that the diffusion operators defined on each modality are relatively similar. This assumption is empirically supported by the results in Sec. 5, as well as in recent works (Huh et al., 2024; Fan et al., 2025). Then, universality is enabled through the eigenfunction preservation properties of the similar diffusion operators. Namely, the eigenfunctions of these operators will be universal, in the sense of modality-invariance (see Sec. 1). In practice, the ability to learn the diffusion operators’ eigenfunctions is obtained via SpectralNet Shaham et al. (2018): while trained to compute the eigenvectors of the random walk matrix of its training data, being a generalizable parametric map makes it a practical means to compute the eigenfunctions of the diffusion operator, viewing the eigenvectors as a discretization of the eigenfunctions (Belkin and Niyogi, 2006; Shi, 2015). Crucially, we train SpectralNet on unimodal data only; hence, no paired data is needed to learn the Laplacian eigenfunctions, i.e., our universal embedding functions.
@@
SUE consists of three steps: SE, CCA and MMD. First, it maps each pre-trained unimodal embedding space into its corresponding eigenspace, to retrieve the global structure of each modality Belkin and Niyogi (2003); Nadler et al. (2006); Singer and Coifman (2008). Using SpectralNet Shaham et al. (2018), this is done parametrically, allowing generalization to test data. Noteworthy, SE is not unique, as eigenvalues with multiplicity $p$ can yield any basis spanning the $p$-dimensional eigenspace and even single eigenvectors may differ by sign.
To resolve the SE ambiguity, and provide additional linear alignment, we use CCA on a minimal number of paired samples. However, as the CCA purposefully considers a limited number of samples, and the SEs differ by more than an orthogonal transformation, we strengthen the cross-modal alignment using a Maximum Mean Discrepancy (MMD) residual network Shaham et al. (2017). This kind of network architecture was originally designed for batch-effect removal, by minimizing the empirical MMD value of two distributions. Namely, we view the two low-dimensional representations as similar distributions and learn a (close to identity) non-linear shift to align the distributions. The MMD serves as the last step to fine-tune the alignment. Notably, MMD loss does not require paired data, which enables the utilization of the full unpaired dataset.

### v2

@@ Abstract
Learning shared representations is a primary area of multimodal representation learning. The current approaches to achieve a shared embedding space rely heavily on paired samples from each modality, which are significantly harder to obtain than unpaired ones. In this work, we demonstrate that shared representations can be learned almost exclusively from unpaired data. Our arguments are grounded in the spectral embeddings of the random walk matrices constructed independently from each unimodal representation. Empirical results in computer vision and natural language processing domains support its potential, revealing the effectiveness of unpaired data in capturing meaningful cross-modal relations, demonstrating high capabilities in retrieval tasks, generation, arithmetics, zero-shot, and cross-domain classification. This work, to the best of our knowledge, is the first to demonstrate these capabilities almost exclusively from unpaired samples, giving rise to a cross-modal embedding that could be viewed as universal, i.e., independent of the specific modalities of the data. Our project page: https://shaham-lab.github.io/SUE_page.
@@ 1 Introduction
The great success of unimodal models [19, 51, 62, 11, 9, 2] in the last decade, and the increasing demand for multimodal applications, have shifted large research attention into the cross-modal domain. Multimodal models present impressive performance on various cross-modal tasks, including image-text [60, 46, 47], speech-text [22], video-text [80] and medical-image-text [84], to name a few.
A primary task in multimodal representation learning is the learning of shared representations, i.e., representation spaces to which instances from different modalities can be mapped, and in which such multimodal instances can be compared. Currently, models learning shared representations typically rely on vast amounts of paired data for training. For example, CLIP [60] was trained on 400 million pairs of images and their captions (i.e., corresponding texts). This kind of supervision (i.e., pairing) is often costly, difficult, and might be even impossible to obtain in many domains and applications. In medical domains, for instance, obtaining a large number of samples is often challenging, as it typically depends on expert annotations. This supervision becomes both costly and time-intensive, or requires rare and specialized tests [32]. Unlike paired data, unpaired data is significantly more accessible and available.
Figure 1: Empirical demonstration of universality. (a) Distances between corresponding random walks on image and text graphs from MSCOCO, compared to distances to randomly shuffled (non-matching) walks. Although constructed independently from unimodal features, corresponding walks exhibit significantly greater similarity. (b) Distances between paired and unpaired points in the shared space of aligned 2D spectral embeddings (SEs). Paired points are consistently closer, indicating that the independently learned SEs capture analogous structure across modalities (see App. A.7).
Figure 2: Almost exclusively unpaired image retrieval. Retrieved images by SUE for custom captions on MSCOCO, trained with 100 pairs and 10k non-pairs. Despite minimal paired data, the results semantically align closely with text queries.
@@
Specifically, modern pre-trained unimodal foundation models have a proven ability to represent semantics [36]. For example, two given images have close embeddings if their semantic meaning is similar, and far otherwise. These similarities can be captured by a random walk process on the samples’ representations, suggesting that a random walk process defined on such unimodal representations should largely correspond to semantic similarity. Therefore, we can expect random walks defined on different unimodal representations that capture semantics well to be highly similar (see Fig. 1). This assumption is supported by substantial empirical evidence from recent works [36, 25].
Random walk processes are finite analogs of diffusion operators. Thereby, the similarity of random walks that are constructed from different, modality-dependent representations implies that the eigenfunctions of the corresponding diffusion operators will have universality properties (i.e., modality-invariance) [13]. Therefore, constructing a spectral embedding (SE) based on the leading eigenvectors of random walks, which are viewed as discrete approximations of the leading eigenfunctions of diffusion operators [5, 69], enables us to take advantage of this concept even in the absence of paired samples. That is, these eigenfunctions can be learned independently from each modality using only unpaired data.
We demonstrate our arguments via a pipeline we call SUE: Spectral Universal Embedding, designed to uncover a universal embedding in a weakly-paired scenario, i.e., when large amounts of unpaired data are available alongside only a very small number of paired samples. To the best of our knowledge, this work is the first to learn shared representations from almost exclusively unpaired data. SUE consists of three steps (see Fig. 3): (i) SE, to extract universal features, (ii) Canonical Correlation Analysis (CCA) on very few samples, to align those features, and (iii) MMD-net, to provide additional non-linear alignment. We demonstrate SUE on vision and language domains and show that with only a small number of paired examples, unpaired data can be leveraged to learn rich multimodal representations. For instance, on the MSCOCO dataset, SUE improves retrieval performance by over 250% on average compared to contrastive learning trained with the same number of pairs. Fig. 2 demonstrates SUE’s ability to learn cross-modal properties with almost no paired data.
@@ 2 Related Work
Multimodal models have recently seen great improvements in rich-data modalities, such as text with image [60, 46, 47], speech [22] and video [80]. Current state-of-the-art multimodal models rely on contrastive learning between corresponding pairs of unimodal representations. While these methods yield high-quality results, they require extremely large amounts of paired data, limiting their applicability to arbitrary domains. Alternative methods, such as CSA [48], slightly reduce the number of pairs required but still operate exclusively on paired data. Unlike these, the universal embedding offers an integration of any two domains, uncovering a shared space primarily from unpaired samples, complemented by only a small number of pairs. Moreover, following recent research on the modality gap in multimodal learning [49, 23, 65], contrastive-based methods fail to be universal due to modalities residing on different manifolds. In contrast, a universal embedding should not exhibit this gap (see App. B for details).
@@
Unpaired samples have been used to replace certain applications of multimodal models without constructing shared representations, including medical image segmentation [82, 75], image-to-speech translation [52], image-to-image translation [87, 88, 12], bilingual word embeddings [1], and domain confusion [26, 27, 72, 85]. Other methods incorporate unpaired data to complement additional signals, such as large paired datasets [56, 71, 18, 28, 78, 83] or classification labels [74, 79, 45]. In contrast, SUE uncovers a latent universal space between two unimodal domains that share semantic content, using unpaired data and only a small number of pairs.
Several works explore shared representation learning from limited to no supervision. Hoshen and Wolf [34] proposed an unpaired variant of CCA, but as noted in their work, the method is unstable and requires many runs to yield satisfactory results, limiting its practicality. MACK [35], while targeting image-text alignment, depends on a segmentation-to-text model trained with paired data, limiting its scope. Closer to our direction are [10, 54], which align modalities by referencing a set of paired examples. However, they overlook local structure in the unimodal embedding space - an important cue, since non-immediate neighbors are often uninformative in high-dimensional settings. In contrast, SUE leverages local unimodal neighborhoods and a few weak pairs to construct a unified cross-modal representation.
@@
SE is known for its global structure preservation properties of unimodal manifolds [15, 55, 73]. One perspective, for instance, is that the SE is the space in which the Euclidean distances represent the diffusion distances on the original manifold [14]. The diffusion distance is a manifold-dependent metric, offering advantages over traditional metrics (e.g., Euclidean, cosine), especially in high-dimensional spaces [15]. Recent works investigated the global structure preservation abilities of SE in the cross-modal scenario [24, 44, 81]. Specifically, they showed that joint SE extracts meaningful representations from each modality. The universal embedding is inspired by these findings, and the analysis we provide extends this idea. Particularly, we show that independent SEs capture similar semantic properties across different modalities, such as images and texts.
@@ 3 Preliminaries
SE [4, 14] is a popular method, used across various domains [53, 16, 86, 3, 21, 42, 43, 68, 8, 89]. In particular, let $\mathcal{X}\subseteq\mathbb{R}^{d}$ denote a collection of data points sampled from a manifold $\mathcal{M}$. Let $W\in\mathbb{R}^{n\times n}$ be a positive symmetric graph affinity matrix, with nodes corresponding to $\mathcal{X}$, and let $D$ be the corresponding diagonal degree matrix (i.e., $D_{ii}=\sum_{j=1}^{n}W_{ij}$). The random walk matrix is defined as $P=D^{-1}W$, and the random walk graph Laplacian as L=IPL=I-P, where $I$ is the identity matrix. The eigenvalues of $P$ can be sorted to satisfy $1=\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$ with corresponding eigenvectors $v_{1},v_{2},\dots,v_{n}$ [76].
@@
Given a Riemannian manifold $\mathcal{M}$ the Laplace-Beltrami operator $\Delta$ acts on functions $f:\mathcal{M}\to\mathbb{R}^{k}$ by $\Delta f=\text{div}(\nabla f)$, where div is the divergence and $\nabla$ the gradient. The diffusion operator (heat kernel) at time $t$ is then defined by $P_{t}=e^{-t\Delta}$. For more details please see [14]. Importantly to this work, the Laplacian and random walk matrices converge to the Laplace-Beltrami and diffusion operators, respectively [14, 5, 69].
@@
Recent works proposed parametric SE implementations, such as SpectralNet [67, 6]. These are deep-learning methods, designed to tackle the scalability and generalizability drawbacks of SE. Specifically, they learn a parametric map $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$, which minimizes the Rayleigh quotient
$\mathcal{L}_{\text{spectralnet}}(f)=\frac{1}{n^{2}}\operatorname{\text{Trace}}\big(f(X)^{T}Lf(X)\big)$
@@
MMD [29, 30] is a distance measure between two probability distributions $p,q$. It is defined with respect to a function class $\mathcal{F}$ by
$\operatorname{\text{MMD}}(\mathcal{F},p,q)=\sup_{f\in\mathcal{F}}\big(\mathbb{E}_{x\sim p}f(x)-\mathbb{E}_{x\sim q}f(x)\big)$
@@
$\operatorname{\text{MMD}}^{2}(\mathcal{F},p,q)=\mathbb{E}_{x,x^{\prime}\sim p}k(x,x^{\prime})-2\mathbb{E}_{x\sim p,y\sim q}k(x,y)+\mathbb{E}_{y,y^{\prime}\sim q}k(y,y^{\prime})$
@@ 4.1 Rationale
We formalize our assumption as follows. Let $\mathcal{M}$ be a latent, underlying semantic manifold, and let $f,g$ be two transformations, such that $f(\mathcal{M})$ and $g(\mathcal{M})$ represent the two modalities from which we observe samples. There is a body of work specifying conditions under which the spectral properties of $\mathcal{M}$ are preserved under $f,g$. For example, if $f,g$ have bounded distortion and bounded Ricci curvature, the corresponding eigenfunctions of the Laplace-Beltrami operator, as well as of the diffusion operator, on $f(\mathcal{M})$ and $g(\mathcal{M})$ are similar in the $L_{\infty}$ sense [7] (see App. D).
Intuitively, our assumption states that the diffusion operators defined on each modality are relatively similar. This assumption is empirically supported by the results in Sec. 5, as well as in recent works [36, 25]. Then, universality is enabled through the eigenfunction preservation properties of the similar diffusion operators. Namely, the eigenfunctions of these operators will be universal, in the sense of modality-invariance (see Sec. 1). In practice, the ability to learn the diffusion operators’ eigenfunctions is obtained via SpectralNet [67]: while trained to compute the eigenvectors of the random walk matrix of its training data, being a generalizable parametric map makes it a practical means to compute the eigenfunctions of the diffusion operator, viewing the eigenvectors as a discretization of the eigenfunctions [5, 69]. Crucially, we train SpectralNet on unimodal data only; hence, no paired data is needed to learn the Laplacian eigenfunctions, i.e., our universal embedding functions.
@@
SUE consists of three steps: SE, CCA and MMD. First, it maps each pre-trained unimodal embedding space into its corresponding eigenspace, to retrieve the global structure of each modality [4, 55, 70]. Using SpectralNet [67], this is done parametrically, allowing generalization to test data. Noteworthy, SE is not unique, as eigenvalues with multiplicity $p$ can yield any basis spanning the $p$-dimensional eigenspace and even single eigenvectors may differ by sign.
To resolve the SE ambiguity, and provide additional linear alignment, we use CCA on a minimal number of paired samples. However, as the CCA purposefully considers a limited number of samples, and the SEs differ by more than an orthogonal transformation, we strengthen the cross-modal alignment using a Maximum Mean Discrepancy (MMD) residual network [66]. This kind of network architecture was originally designed for batch-effect removal, by minimizing the empirical MMD value of two distributions. Namely, we view the two low-dimensional representations as similar distributions and learn a (close to identity) non-linear shift to align the distributions. The MMD serves as the last step to fine-tune the alignment. Notably, MMD loss does not require paired data, which enables the utilization of the full unpaired dataset.